SubjectCover

Question 12 of 18 in Handling Data 2

Handling Data 2

2 marks · Normal

A bag contains 2 red balls and 3 blue balls. One ball is chosen at random, replaced, and another is chosen.

What is the probability that both balls are red?

Work this out on paper, then compare with the answer and worked solution below. With a free account you can submit written answers for marking.

Show answer

4/25 (±0.01)

Also accept: 0.16; 4/25; 0.16 (2 d.p.)

Common mistakes: Using combinations, not product rule

Show mark scheme and worked solution

Marking points

M1 · 1 mark · Identifies probability for one red ball: 2/5

A1 · 1 mark · Correctly squares to get (2/5) × (2/5) = 4/25 or 0.16

Worked solution

Probability red on first pick: 2/5

Probability on second pick (after replacing): 2/5

Probability both red: 2/5 × 2/5 = 4/25 = 0.16

Report a problem with this question
What's wrong?